OpenStudy (anonymous):

what is the length of the side x? if necessary round your answer to two decimal places the hypotenuse is 19 and the leg is 12

OpenStudy (callisto):

Is it a right angle triangle?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

So, by Pyth Thm, \[19^2 = 12^2 + x^2\]\[x= \sqrt{19^2-12^2}\]Can you do it?

OpenStudy (anonymous):

im not that smart lol

OpenStudy (callisto):

Use your calculator?

OpenStudy (anonymous):

i only have a standard calculator it doesnt have all that

OpenStudy (callisto):

First enter '19^2 - 12^2' Then take square root of the answer

OpenStudy (callisto):

19^2 = 361 12^2 = 144

OpenStudy (anonymous):

so i find the square root of 505 on the first one?

OpenStudy (callisto):

Nope!!

OpenStudy (anonymous):

ok well what do i find the square root of

OpenStudy (callisto):

first find the difference between 19^2 and 12^2, can you?

OpenStudy (anonymous):

19 is 361 ?

OpenStudy (callisto):

19^2 is 361

OpenStudy (anonymous):

12 is 144

OpenStudy (callisto):

and the difference is?

OpenStudy (anonymous):

i dont know can you explain how to get it

OpenStudy (anonymous):

217

OpenStudy (anonymous):

i think the answer is 14.73

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

thank you :)

OpenStudy (callisto):

Welcome!